In a recent poll, 600 people were asked if they liked dogs, and 14% said they did. Find the margin of error of this poll, at the 90% confidence level.
Answer #1

p= 0.14

q= 1-p=1-0.14=0.86

n= 600

alpha=0.1 then Z(alpha/2)=1.645

Margin of error ,

p*q MOE Z(al pha/2)

0.14 *0.86 1.645/ 600

=0.02330

if you have any doubt ask in comment give thumbs up if you like wor.